What is the vertex of the parabola defined by---
y=-2(x+3)^2+5
The answer choices are; A. (-3,5) B. (-3,-5) C. (3,5) D. (3,-5)

Found 2 solutions by richwmiller, Edwin McCravy:
Answer by richwmiller(17219)

no similar problems

Answer by Edwin McCravy(20055)

Learn the rule:

y = a(x - h)² + k 

has vertex (h, k). 

Notice that h has the OPPOSITE sign from the sign just before it, 
and that k has the SAME sign that is before it.

What is the vertex of the parabola defined by---
y = -2(x + 3)² + 5
 
The 3 has a + before it so we use the OPPOSITE sign -.
The 5 has a + before it so we use the SAME sign +.
So the vertex is (-3,5)

What is the vertex of the parabola defined by---
y = (x + 4)² - 6
The 4 has a + before it so we use the OPPOSITE sign -.
The 6 has a - before it so we use the SAME sign -.
So the vertex is (-4,-6)

Edwin